Can you help with letter b? For letter a I got 93 bottles idk if thats correct but not sure how to do b? And I know S= R(l+T) where l is lead time which is the 3 days

The manager of store Z asks you to help in determining the optimal ordering policy for Truck Stop Honey. The demand for this beer is 35 bottles per day. The store pays a unit cost of $3.50 per bottle and a fixed ordering cost of $130 per order. The annual unit holding cost of inventory is estimated to be 30% of the unit cost of each bottle. The store operates 280 days a year.

a. Determine the Economic Order Quantity (EOQ).

b. Compute the reorder point (ROP) assuming that the order lead time is 3 days.
Answer #1

A. EOQ = SQRT(2 * DEMAND * ORDERING COST / HOLDING COST)

Annual demand = 35 * 280 = 9800

Holding cost = 30% of $3.5

EOQ = SQRT(2 * 9800 * 130 / 1.05) = 1557.78

B. ROP = daily demand * Lead time = 35 * 3 = 105

In EOQ question, when safety stock is not mentioned, you can calculate ROP by using the above formula.

Comment if you have any problem.

Do give a positive rating if you think this answer helped